OpenStudy (anonymous):

Janis and Susan went to a clearance sale at a clothing store. All blouses were selling for the same price and all skirts were selling for the same price. Janis bought 4 blouses and 3 skirts for a total of $68. Susan bought 6 blouses and 3 skirts for a total of $84. What was the cost of each blouse and each skirt? A. blouse: $12; skirt: $12 B. blouse: $12; skirt: $8 C. blouse $10; skirt $14 D. blouse: $8; skirt $12 <--- @matt101 @Hero @tHe_FiZiCx99

OpenStudy (anonymous):

@mathslover @matt101 @Hero @tHe_FiZiCx99 @jim_thompson5910

OpenStudy (anteater):

They don't tell you how much each skirt costs or how much each blouse costs, so we can use variables to stand for those amounts until we figure out what they are.

OpenStudy (anteater):

So, you could use, for example, x to stand for the cost of a skirt and y to stand for the cost of one blouse.

OpenStudy (anonymous):

It as pretty easy to figure out... I just wanted to know if I was right for sure. :) @ParthKohli @zepdrix @Luigi0210 @wio @iambatman @iPwnBunnies

OpenStudy (anteater):

Oh, ok. :)!

OpenStudy (anteater):

Just a moment, please. 8X4 + 3 X12 = 32 + 36 = $68

OpenStudy (anonymous):

8x4=32 12x3=36 36+32=68 8x6=48 12x3=36 48+36=84

OpenStudy (anteater):

and 6 X8 + 3X12 = 48 + 36 = $84 So, yes :)

OpenStudy (anteater):

Sorry, I didn't see that you had marked the answer and just wanted to check it. :)

OpenStudy (anteater):

And yeah, I got the same thing.

OpenStudy (anteater):

So ... good :D

OpenStudy (anonymous):

KK, I'm just going to see if I can get one more person to agree. Better safe than sorry. ;P @iambatman

OpenStudy (anonymous):

@iambatman

OpenStudy (anteater):

You are pretty safe! :D But caution is never a bad thing in math. :)

OpenStudy (anteater):

Have a good evening!

OpenStudy (anonymous):

Ivan has nickels and dimes in his pocket. The total number of nickels and dimes is 16. The total value of the coins is $1.25. How many of each coin does he have? A. 6 nickels and 10 dimes B. 7 nickels and 9 dimes <----- C. 8 nickels and 8 dimes D. 9 nickels and 7 dimes 9 dimes = 90 cents 7 nickels = 35 cents 90+35 cents = 1.25

OpenStudy (anonymous):

@iambatman Are both of those right?

OpenStudy (anteater):

7 X .05 + 9 X .10 = .35 + .90 = $1.25 Yes :)

OpenStudy (anteater):

G'night.

OpenStudy (anonymous):

@anteater Can you check two more? :)

OpenStudy (anteater):

Sure!

OpenStudy (anonymous):

Jeff bought 5 oranges and 5 apples for a total of $7.70. Gene bought 5 oranges and 3 apples for a total of $6.40. What is the cost of each orange and each apple? A. orange: $0.65; apple: $0.89 <--- B. orange: $0.75; apple: $0.79 C. orange: $0.79; apple: $0.75 D. orange: $0.89; apple: $0.65

OpenStudy (anteater):

5 X .65 + 5 X .89 = 3.25 + 4.45 = $7.70 and 5 X .65 + 3 X .89 = 3.25 + 2.67 = $5.92 ... That didn't work out, so let me double check to make sure!

OpenStudy (anteater):

OOps

OpenStudy (anteater):

I mixed up apples and oranges on the second one

OpenStudy (anonymous):

Gio drew an angle on a coordinate grid. One side of the angle had the equation 2x − 3y = 0, and the other side had the equation 2x + 9y = 24. What were the coordinates of the vertex of the angle Gio drew? A. (2, 3) B. (2, −3) C. (3, 2) <---- D. (−3, 2)

OpenStudy (anteater):

3 X .65 + 5 X .89 = $6.40 So, yes :D

OpenStudy (anonymous):

What about the next one? :)

OpenStudy (anteater):

2(3) - 3(2) = 0 and 2(3) + 9(2) = 6 + 18 = 24 so yes :D

OpenStudy (anteater):

Now, let me think on that last one just to make sure I didn't overlook something

OpenStudy (anteater):

That is an oddly worded problem, actually! But, I believe they are simply wanting you to find the solution to the system of linear equations, so yes (3,2) would work.

OpenStudy (anteater):

I see what they are saying. It would have been better worded by saying that two lines meet to form an angle. :)

OpenStudy (anteater):

And the vertex of the angle would be the point where those lines intersect, which is (3,2), so yes you are correct. :)

OpenStudy (anteater):

Got any others?

OpenStudy (anteater):

If not that's ok ... anteater will go eat chocolate and knit. :)

OpenStudy (anteater):

Have a fine evening! :D